OpenStudy (anonymous):

A students desk at home is rectangular in shape and measures 3 feet wide by 5 feet long. Find the perimeter of the desk in centimeters. Then convert the answer to decimeters. Please help :)

OpenStudy (anonymous):

3 ft X 5 ft; so 3(2) + 5(2) = 6 + 10 = 16ft^2. So there are 12 in. in a foot, so that's 192in. There are 2.54 centimeters in an inch, so then you can say that there are 487.68 centimeters in 192 inches. Then to decimeters; 10 are in 1 centimeter, so that gives you 4876.8 decimeters, right? :)

OpenStudy (anonymous):

WAIt sorry no no no no there are 0.1 decimeters in a centimeter, SORRY. Its 487.68 * 0.1 = 48.768dm I think

OpenStudy (anonymous):

I know the answer is 141.4 yards.....but I'm lost on how to figure it out.

OpenStudy (radar):

The problem it wants the answer in centimeters, then that converted to decimeters. Where do the 141.4 yds figure?

OpenStudy (radar):

I suspect you are looking at an answer for a different question.

OpenStudy (anonymous):

okay so sorry 16 foot

OpenStudy (radar):

O.K yes the perimeter is 16 foot, but that is an intermediate step, now you must convert feet to centimeters.

OpenStudy (radar):

I know 4 feet = 121.92, so 16 ft = 4 times 121.92=487.68 cm

OpenStudy (anonymous):

how is the 16 feet figured?

OpenStudy (radar):

1 cm = .1 dm, so 487.68 (.1)= 48.768

OpenStudy (radar):

Perimeter is the sum of the sides of the desk, 5+5+3+3=16 ft.

OpenStudy (anonymous):

yay! I understand thanks so much.....

OpenStudy (radar):

Great. Good luck.